# A fluid–peridynamic structure model of deformation and damage of microchannels

### v1

@@ 1 Introduction
In classical continuum mechanics, each material point is assumed to interact solely with its immediate neighbors [1, 2]. This local interaction paradigm is typically captured through partial differential equations (PDEs) that utilize spatial derivatives. While this framework has proven successful in modeling a wide range of mechanical phenomena, it encounters significant challenges in at least two important scenarios. First, classical continuum theory struggles with discontinuities, where the field variables may be undefined or exhibit sharp gradients—such as during fracture, the formation of cracks, and sudden material separations [3]. Because the classical theory relies on being able to compute spatial derivatives of continuous fields, it lacks a straightforward means to handle abrupt damage within the material. Second, and related to the previous point, certain multiphysics phenomena are intrinsically nonlocal [4], requiring knowledge of stresses, strains, or more generally the state of the body in the neighborhood around the material point under question (not just the nearest neighbor material points). Nonlocality can also be thought of as a manifestation of material responses controlled by internal length scales [5].
The theory of peridynamics provides a robust alternative to classical continuum mechanics and avoids many of the limitations imposed by local spatial derivatives [3, 6]. Peridynamics is built on an integral-based equation of motion, making it inherently applicable regardless of whether discontinuities are present or not. Specifically, in three dimensions, the basic balance law of a peridynamic solid takes the general form
@@
In the formulation (1), each point in the material, often referred to as the “center particle” $\mathbf{x}$, interacts with every other point $\mathbf{x}^{\prime}$ within a finite region (neighborhood), $\mathcal{H}(\mathbf{x})$, surrounding $\mathbf{x}$. This finite region is known as the horizon and is characterized by the horizon size $\delta$, an intrinsic length scale. The force functional $\mathbf{f}$ then maps the positions of the center particle and all particles within its horizon to the net force acting at $\mathbf{x}$. By relying on these integral interactions instead of spatial derivatives, peridynamics naturally accommodates cracks, damage, and other material discontinuities without compromising mathematical validity [7, 8]. As a result, the theory of peridyanimics has gained significant attention for modeling fracture initiation, propagation, and dynamic behavior in elastic materials [9, 6, 10, 11, 12].
The interaction between a two-dimensional fluid flow and a compliant structure governs the performance and reliability of systems that span geophysical, biomedical, and micro-engineering scales. In the Earth’s mantle, for example, channelized magma flow can weaken overlying lithospheric plates and initiate large-scale failure processes such as earthquakes and volcanic eruptions [13, 14], leading to the creation of laccolith formations [15, 16, 17, 18]. At the opposite length scale, organ-on-a-chip devices rely on a microfluidic perfusion layer to supply nutrients to delicate cell cultures; excessive shear or pressure in the flow damages the soft cell layer on the channel walls and compromises tissue viability [19, 20]. Related concerns arise in microfluidic devices fabricated from polydimethylsiloxane (PDMS), where high-throughput flows can cause the thin PDMS ceilings of channels to fracture [21].
@@
This study couples the peridynamic formulation for a thin, compliant channel wall with lubrication theory for the underlying viscous flow. The resulting model captures both the nonlocal solid response and the fluid–structure interaction that drives deformation and damage. Section 2 formulates the physical problem. To this end, this study is organized as follows. Section 3 develops the dimensionless governing equations, introduces the key dimensionless groups—the horizon size $\Delta$, the Strouhal number $St$, the compliance number $\beta$, and the Reynolds number ReRe—and introduces the simulation methodology. Section 4 discusses steady-state (Section 4.1) and transient (Section 4.2) simulations of the coupled fluid–peridynamic structure interaction problem. Section 4.3 presents a dispersion study that traces how the key dimensionless numbers modify the phase velocity and spatial damping rate in this system. Then, Section 4.4 discusses the fracture risk comparison between dynamic loading and an equivalent static-pressure loading, showing when and how the channel’s ceiling may fail. Section 5 concludes our study and discusses avenues for future work.
@@ 2 Problem formulation
Following Inamdar et al. [40], we seek to address the fluid–structure interaction problem in which a 2D channel with a top elastic wall of length $\ell$ is clamped at both ends. The top wall is modeled as a beam, with transverse (vertical) displacement $u(x,t)$. The channel’s undeformed height is $h_{0}$. Thus, its deformed height is given by $h=h_{0}+u$.
@@ 3.1 Peridynamic beam theory
The transverse displacements of a thin beam under bending can be calculated using Euler–Bernoulli beam theory [42]. This simplified version of the linear theory of elasticity provides a framework for analyzing the deflection behavior of beams, specifically in cases of small displacements and strain caused by external loads applied perpendicular to the beam’s longitudinal axis [43].
To motivate the governing equation for the beam, let us first review some basics. Based on the Euler–Bernoulli beam kinematics, the strain tensor $\varepsilon_{ij}$ within the beam can be represented as:
@@
where $\delta$ defines the size of the region of nonlocality, which is the horizon size in the context of peridynamics. It is expected that as $\delta\to 0^{+}$, the peridynamic beam theory should reduce to the classical beam theory. Indeed, performing Taylor series expansion of $u$ in $x$, and neglecting higher-order terms, it is straightforward to show that Eq. (5) reduces to Eq. (4) as $\delta\to 0^{+}$. Then, Yang et al. [25, 28, 26] followed an energy‐based methodology, in which the mechanical system’s behavior is described by a Lagrangian. Employing the Euler–Lagrange equations to extremize the action, they obtained the equation of motion of the transverse displacement of a loaded peridynamic beam based on the strain-energy density from Eq. (5). Specifically,
@@
In our problem, both ends of the beam are assumed to be clamped (see Fig. 1), thereby restricting two degrees of freedom at each boundary: the displacement $u$ and the rotation (slope, $\partial u/\partial x$) [43]. Mathematically, these constraints are imposed by prescribing zero Dirichlet and zero Neumann boundary conditions (BCs) on the displacement $u(x,t)$ at $x=0$ and $x=\ell$, specifically:
$u(0,t)$ $=0,\qquad$ $\frac{\partial u}{\partial x}(0,t)$ $=0,$ (7a)
$u(\ell,t)$ $=0,\qquad$ $\frac{\partial u}{\partial x}(\ell,t)$ $=0.$ (7b)
Here, the zero Dirichlet conditions fix the beam’s displacement at the two ends, while the zero first-derivative conditions eliminate any rotation, accurately reflecting the physical scenario of a beam rigidly clamped at both supports.
@@ 3.2 Lubrication theory incorporating flow inertia
We proceed by adopting the nondimensionalization of Inamdar et al. [40] based on lubrication theory. To this end, we introduce the following dimensionless (capital) variables:
@@
A no-slip BC is imposed on both the top and bottom walls of the channel. Since the top wall is in motion, the no-slip condition at the top wall also implies a kinematic (boundary) condition at $Y=H(X,T)$ to match the vertical velocities of the flow and structure [33]. Enforcing the small-slope approximation of beam theory ($\partial H/\partial X\ll 1$), the kinematic condition simplifies [40] to:
@@
Next, we also integrate the $X$-momentum equation (11b) from $Y=0$ to $Y=H(X,T)$. To do so, first, it is transformed into conservative form, then the horizontal velocity is assumed to follow a Hagen–Poiseuille (parabolic) flow profile: $V_{X}={6Y(H-Y)Q}/{H^{3}}$, which is a von Kármán–Pohlhausen-type of closure [40, 44, 45]. This closure is asymptotically consistent and valid at low Womersley number (i.e., for low flow oscillation frequency) [46, 47], which is consistent with our time-independent flow BCs given in Eq. (9). The pressure gradient is independent of $Y$ thanks to the $Y$-momentum equation (11c). Performing these (lengthy) calculations, we arrive at the 1D form of the axial momentum equation:
@@ 3.3 Summary of the model’s governing equations
The Strouhal number $St$ quantifies the beam’s unsteady inertia; a smaller $St$ corresponds to a faster loading response. Meanwhile, $\beta$ quantifies the fluid–structure interaction, combining the principal parameters that govern the fluid flow (viscosity and flow rate) and the structural deformation (bending rigidity); a smaller $\beta$ corresponds to weaker coupling between the flow and structure. Finally, the dimensionless BCs from Eq. (7) for the peridynamic beam governed by Eq. (16), corresponding to clamping at the ends, become:
$H(0,T)$ $=1,\qquad$ $\frac{\partial H}{\partial X}(0,T)$ $=0,$ (19a)
$H(1,T)$ $=1,\qquad$ $\frac{\partial H}{\partial X}(1,T)$ $=0.$ (19b)
@@ 3.4 Solution approach and parameter space
The governing equations are discretized in space and time to solve for the dynamic responses of the system. The 1D axial domain is uniformly discretized, and the time integration is performed using the Newmark-$\beta$ time-stepping method [48]. The technical details of the numerical methodology for the proposed 1D fluid–peridynamic structure model are given in A. Briefly, we use a segregated iterative approach following [40, 41]. We first solve the beam’s motion equation (16) given the pressure $P$ to obtain the deformed channel height $H$ and its time derivative ${\partial H}/{\partial T}$. Then, the flow rate $Q$ is found from the continuity Eq. (14) by integrating ${\partial H}/{\partial T}$ over $X$. Finally, the pressure is computed from the axial momentum Eq. (15) using second-order backward differences in time combined with central differences in space. These calculations are iteratively repeated until the solution converges within each time step.
@@ 4.1 Steady state and validation
As the beam first responds to the flow, it undergoes a brief oscillatory transient before settling to a steady configuration. To shorten the transient time window in this steady-state benchmark, we reduced the Strouhal number $St$, thereby diminishing inertial effects and accelerating convergence. Steady state is considered achieved when the relative change in the displacement field between successive time steps, measured using the $L^{2}$ norm, falls below $10^{-5}$. Once the steady state is achieved, we examine the deformed shape of the beam and the corresponding pressure distribution along its length. We focus on the case of $\beta=556$ and $Re=0.5$ from [40].
Figure 2: Steady state (a) beam deformations $H(X)$ and (b) pressure load $P(X)$ along the beam computed using classical continuum mechanics by Inamdar et al. [40] (solid curves) and the fluid–peridynamic structure solver developed in this work (dashed curves), for three horizon sizes: $\Delta=1/100$ and $\Delta=1/20$. The simulation parameters are $\beta=556$ and $Re=0.5$.
Fig. 2 compares the peridynamics-based results with those obtained using classical continuum mechanics in [40]. As seen from the figure, both the beam deformation and the pressure distribution along the beam are in close agreement with the solution from the literature. Notably, reducing the horizon size yields an almost identical solution to the classical counterpart, indicating that the peridynamic theory is converging towards the classical one as the horizon size is decreased, as expected.
@@ 4.2 Nonlinear dynamics and effect of the horizon
Next, we discuss the dynamic simulation, and a sample result is shown in Fig. 3. We observe that, in general, the flow-induced deformation manifests as a bulge that originates near the flow inlet and migrates toward the middle of the channel. The main bulge-like deformation is accompanied by small wave-like perturbations. As the bulge travels downstream, the maximum height of the channel gradually increases. However, once the bulge reaches the middle region, the peak’s height begins to fluctuate up and down. The bulge begins to oscillate laterally, shifting slightly from left to right around the centerline of the channel. These oscillations eventually subside as the system approaches a steady-state configuration, as discussed in Sec. 4.1. The frequency and amplitude of the vertical and horizontal oscillations are strongly influenced by the dimensionless parameters $St$, $\beta$, and $Re$, reflecting the nonlinear interplay between beam inertial, fluid–structure coupling, and viscous forces in this problem.
Figure 3: Nonlinear dynamics of the coupled problem illustrated by the time evolution of the channel wall height, $H(X,T)$. Darker curves represent later times, while lighter curves correspond to earlier times. The red arrows indicate the direction of temporal evolution of the channel height. The simulation parameters are $St=10$, $\beta=20$, $Re=0.5$, and $\Delta=1/120$.
The unsteady evolution is qualitatively similar to that in [40] under the classical continuum theory, as expected. To get a sense of the peridynamic aspect of the present problem, we next determine the influence of the dimensionless horizon length $\Delta$ on the dynamics. This influence is most easily observed at early times. Fig. 4 shows the deformed channel height recorded at an early time during the simulation ($T=0.02$), for four different horizon sizes. The remaining parameters are fixed at $St=10$, $\beta=10^{5}$, and $Re=10^{-5}$ for these simulations. From Fig. 4, it is evident that the smaller oscillations to the right of the primary bulge undergo stronger damping when the horizon size is larger (e.g., $\Delta=1/30$ and $1/24$ compared to $\Delta=1/20$ and $1/12$). For $\Delta=1/12$, these oscillations have nearly flattened, indicating more significant energy dissipation for stronger nonlocality and more effective dissipation of vibrational energy in the system.
Figure 4: Effect of the horizon size $\Delta$ on the damping behavior of the system at an early time of $T=0.02$, showing the stronger wave damping as $\Delta$ increases. The simulation parameters are $St=10$, $\beta=10^{5}$, and $Re=10^{-5}$.

### v2

@@ 1 Introduction
In classical continuum mechanics, each material point is assumed to interact solely with its immediate neighbors [49, 15]. This local interaction paradigm is typically captured through partial differential equations (PDEs) that utilize spatial derivatives. While this framework has proven successful in modeling a wide range of mechanical phenomena, it encounters significant challenges in at least two important scenarios. First, classical continuum theory struggles with discontinuities, where the field variables may be undefined or exhibit sharp gradients—such as during fracture, crack formation, and sudden material failure or separation [39]. Because the classical theory relies on being able to compute spatial derivatives of continuous fields, it lacks a straightforward means to handle abrupt damage within the material. Second, and related to the previous point, certain multiphysics phenomena are intrinsically nonlocal [53], requiring knowledge of stresses, strains, or more generally the state of the body in the neighborhood around the material point under question (not just the nearest neighbor material points). Nonlocality can also be thought of as a manifestation of material responses controlled by internal length scales [1].
The theory of peridynamics provides a robust alternative to classical continuum mechanics and avoids many of the limitations imposed by local spatial derivatives [39, 37]. Peridynamics is based on an integral equation of motion, making it inherently applicable regardless of whether discontinuities are present. Specifically, in three dimensions, the basic balance law of a peridynamic solid takes the general form
@@
In the formulation (1), each point in the material, often referred to as the “center particle” $\mathbf{x}$, interacts with every other point $\mathbf{x}^{\prime}$ within a finite region (neighborhood), $\mathcal{H}(\mathbf{x})$, surrounding $\mathbf{x}$. This finite region is known as the horizon and is characterized by the horizon size $\delta$, an intrinsic length scale. The force functional $\mathbf{f}$ then maps the positions of the center particle and all particles within its horizon to the net force acting at $\mathbf{x}$. By relying on these integral interactions instead of spatial derivatives, peridynamics naturally accommodates cracks, damage, and other material discontinuities without compromising its mathematical validity [41, 40]. As a result, the theory of peridynamics has gained significant attention for modeling fracture initiation, propagation, and dynamic behavior in elastic materials [36, 37, 2, 16, 38].
The interaction between a fluid flow and a compliant structure governs the performance and reliability of systems that span geophysical, biomedical, and micro-engineering scales. In the Earth’s mantle, for example, channelized magma flow can weaken overlying lithospheric plates and initiate large-scale failure processes such as earthquakes and volcanic eruptions [24, 42], leading to the creation of laccolith formations [26, 5, 4, 48]. At the opposite length scale, organ-on-a-chip devices rely on a microfluidic perfusion layer to supply nutrients to delicate cell cultures; excessive shear or pressure in the flow damages the soft cell layer on the channel walls and compromises tissue viability [28, 14]. Related concerns arise in microfluidic devices fabricated from polydimethylsiloxane (PDMS), where high-throughput flows can cause the thin PDMS ceilings of channels to fracture [62].
@@
This study couples a peridynamic formulation for a thin, compliant channel wall with a lubrication description of the underlying viscous flow. The resulting model captures both the nonlocal solid response and the fluid–structure interaction that drives deformation and damage. To this end, this study is organized as follows. Sec. 2 formulates the physical problem. Sec. 3 develops the dimensionless governing equations, introduces the key dimensionless groups—the horizon size $\Delta$, the Strouhal number $St$, the compliance number $\beta$, and the Reynolds number ReRe—and introduces the simulation methodology. Sec. 4 discusses steady-state (Sec. 4.1) and transient (Sec. 4.2) simulations of the coupled fluid–peridynamic structure interaction problem. Sec. 4.3 presents a dispersion study that traces how the key dimensionless numbers modify the phase velocity and spatial damping rate in this system. Then, Sec. 4.4 compares the fracture ‘risk’ between dynamic loading and an equivalent static-pressure loading, showing when and how the channel’s ceiling may fail. Sec. 5 discusses some of the advantages, as well as limitations, of our modeling approach, while Sec. 6 concludes our study and discusses avenues for future work.
@@ 2 Problem formulation
Following Inamdar et al. [20], we seek to address the fluid–structure interaction problem in which a two-dimensional (2D) channel with a top elastic wall of length $\ell$ is clamped at both ends. The top wall is modeled as a beam, with transverse (vertical) displacement $u(x,t)$. The channel’s undeformed height is $h_{0}$. Thus, its deformed height is $h=h_{0}+u$.
@@ 3.1 Peridynamic beam theory
The transverse displacements of a thin beam under bending can be calculated using Euler–Bernoulli beam theory [18]. This simplified version of the linear theory of elasticity provides a framework for analyzing the deflection behavior of beams, when the external loads applied perpendicular to the beam’s longitudinal axis cause only small displacements and strains [34].
To motivate the governing equation for the beam, let us first review some basics. Based on the Euler–Bernoulli beam kinematics, the strain tensor $\varepsilon_{ij}$ within the beam can be represented as
@@
where $\delta$ defines the size of the region of nonlocality, which is the horizon size in the context of peridynamics. It is expected that as $\delta\to 0^{+}$, the peridynamic beam theory should reduce to the classical beam theory. Indeed, performing Taylor series expansion of $u$ in $x$, and neglecting higher-order terms, it is straightforward to show that Eq. (5) reduces to Eq. (4) as $\delta\to 0^{+}$. Then, Yang et al. [56, 60, 58] followed an energy‐based methodology, in which the mechanical system’s behavior is described by a Lagrangian. Employing the Euler–Lagrange equations to extremize the action, they obtained the equation of motion of the transverse displacement of a loaded peridynamic beam based on the strain energy density from Eq. (5). Specifically,
@@
In our problem, both ends of the beam are assumed to be clamped (see Fig. 1), thereby restricting two degrees of freedom at each boundary: the displacement $u$ and the rotation (which is the slope, $\partial u/\partial x$, in the CCM) [34]. Thus, in the classical Euler–Bernoulli beam theory, clamping corresponds to the boundary conditions of $u=0$ and $\partial u/\partial x=0$ at $x=0$ and $x=\ell$. In the PD beam implementation, however, the rotation constraint is imposed using a nonlocal boundary treatment (specifically, by fictitious/ghost material points and symmetry conditions) rather than conditions on spatial derivatives. Therefore, following Yang et al. [56], we formulate the following boundary conditions:
$u(0,t)$ $=0,\quad$ $u(-\zeta,t)$ $=u(\zeta,t),\qquad$ $\zeta\in(0,2\delta],$ (7a)
$u(\ell,t)$ $=0,\quad$ $u(\ell+\zeta,t)$ $=u(\ell-\zeta,t),\qquad$ $\zeta\in(0,2\delta].$ (7b)
Here, the zero Dirichlet conditions fix the beam’s displacement at the two ends. The mirror-image conditions enforce zero rotation in a PD-consistent manner by prescribing the displacement field in a fictitious boundary layer adjacent to each end. We note that there are other PD beam theories that enforce boundary conditions by treating the rotation as an independent kinematic variable [30, 57].
@@ 3.2 Lubrication theory incorporating flow inertia
We proceed by adopting the nondimensionalization of Inamdar et al. [20] based on lubrication theory. To this end, we introduce the following dimensionless variables denoted by capital letters:
@@
A no-slip BC is imposed on both the top and bottom walls of the channel. Since the top wall is in motion, the no-slip condition at the top wall also implies a kinematic (boundary) condition at $Y=H(X,T)$ that matches the vertical velocities of the flow and structure [33]. Enforcing the small-slope approximation inherent to the beam theory ($\partial H/\partial X\ll 1$), the kinematic condition simplifies [20] to:
@@
Next, we also integrate the $X$-momentum equation (11b) from $Y=0$ to $Y=H(X,T)$. To do so, first, it is transformed into conservative form, then the horizontal velocity is assumed to follow a Hagen–Poiseuille (parabolic) flow profile: $V_{X}={6Y(H-Y)Q}/{H^{3}}$, which is a von Kármán–Pohlhausen-type of closure [20, 43, 21]. This closure is asymptotically consistent and valid at low Womersley number (i.e., for low flow oscillation frequency) [52, 63], which is consistent with our time-independent flow BCs given in Eq. (9). The pressure gradient is independent of $Y$ thanks to the $Y$-momentum equation (11c). Performing these (lengthy but standard) calculations, we arrive at the 1D form of the axial momentum equation:
@@ 3.3 Summary of the model’s governing equations
The Strouhal number $St$ quantifies the beam’s unsteady inertia; a smaller $St$ corresponds to a faster loading response. Meanwhile, $\beta$ quantifies the fluid–structure interaction, combining the principal parameters that govern the fluid flow (viscosity and flow rate) and the structural deformation (bending rigidity); a smaller $\beta$ corresponds to weaker coupling between the flow and the structure. Finally, the dimensionless BCs from Eq. (7) for the peridynamic beam governed by Eq. (16), corresponding to clamping at the ends, become:
$H(0,T)$ $=1,\quad$ $H(-Z,T)$ $=H(Z,T),\qquad$ $Z\in(0,2\Delta],$ (19a)
$H(1,T)$ $=1,\quad$ $H(1+Z,T)$ $=H(1-Z,T),\qquad$ $Z\in(0,2\Delta].$ (19b)
@@ 3.4 Solution approach and parameter space
The governing equations are discretized in space and time to determine the system’s dynamic responses. The 1D axial domain is uniformly discretized, and the time integration is performed using the Newmark-$\beta$ time-stepping method [45]. The technical details of the numerical methodology for the proposed 1D fluid–peridynamic structure model are given in A. Briefly, we use a segregated iterative approach following [20, 54]. We first solve the beam’s motion equation (16) given the pressure $P$ to obtain the deformed channel height $H$ and its time derivative ${\partial H}/{\partial T}$. Then, the flow rate $Q$ is found from the continuity Eq. (14) by integrating ${\partial H}/{\partial T}$ over $X$. Finally, the pressure is computed from the axial momentum Eq. (15) using second-order backward differences in time combined with central differences in space. These calculations are iterated until the solution converges at each time step.
@@ 4.1 Steady state and validation
As the beam first responds to the flow, it undergoes a brief oscillatory transient before settling to a steady configuration. To shorten the transient time window in this steady-state benchmark, we reduced the Strouhal number $St$, thereby diminishing inertial effects and accelerating convergence. Steady state is considered achieved when the relative change in the displacement field between successive time steps, measured using the $L^{2}$ norm, falls below $10^{-5}$. Once steady state is achieved, we examine the beam’s deformed shape and the corresponding pressure distribution along its length. We focus on the case of $\beta=556$, and $Re=0.5$ from [20].
Figure 2: Steady state (a) deformed channel wall $H(X)$ and (b) pressure load $P(X)$ along the beam computed using classical continuum mechanics by Inamdar et al. [20] (solid curves) and the fluid–peridynamic structure solver developed in this work (dashed curves), for three horizon sizes: $\Delta=1/100$ and $\Delta=1/20$. The simulation parameters are $\beta=556$ and $Re=0.5$.
Fig. 2 compares our PD-based results with those obtained using CCM in [20]. As seen from the figure, both the beam deformation and the pressure distribution along the beam are in close agreement with the solution from the literature. Notably, reducing the horizon size yields an almost identical solution to the classical counterpart, indicating that the peridynamic theory is converging towards the classical one as the horizon size is decreased, as expected.
@@ 4.2 Nonlinear dynamics and effect of the horizon
Next, we discuss the dynamic simulations. A sample result is shown in Fig. 3. We observe that, in general, the flow-induced deformation manifests as a bulge that originates near the flow inlet and migrates toward the middle of the channel. The main bulge-like deformation is accompanied by small wave-like perturbations. As the bulge travels downstream, the maximum height of the channel gradually increases. However, once the bulge reaches the middle region, the peak’s height begins to fluctuate up and down. The bulge begins to oscillate laterally, shifting slightly from left to right around the centerline of the channel. These oscillations eventually subside as the system approaches the steady state, as discussed in Sec. 4.1. The frequency and amplitude of the vertical and horizontal oscillations are strongly influenced by the dimensionless parameters $St$, $\beta$, and $Re$, reflecting the nonlinear interplay between beam inertia, fluid–structure coupling, and flow inertia in this problem.
Figure 3: Nonlinear dynamics of the coupled problem illustrated by the time evolution of the channel height, $H(X,T)$. Darker curves represent later times, while lighter curves correspond to earlier times. The arrows indicate the direction of temporal evolution of the channel height. The simulation parameters are $St=10$, $\beta=20$, $Re=0.5$, and $\Delta=1/120$.
The unsteady evolution is qualitatively similar to that in [20] under a CCM theory, as expected. To get a sense of the peridynamic aspect of the present problem, we next determine the influence of the dimensionless horizon length $\Delta$ on the dynamics. This influence is most easily observed at early times. Fig. 4 shows the deformed channel height recorded at an early time during the simulation ($T=0.02$), for four different horizon sizes. The remaining parameters are fixed at $St=10$, $\beta=10^{5}$, and $Re=10^{-5}$ for these simulations. From Fig. 4, it is evident that the smaller oscillations to the right of the primary bulge undergo stronger damping when the horizon size is larger (e.g., $\Delta=1/30$ and $1/24$ compared to $\Delta=1/20$ and $1/12$). For $\Delta=1/12$, these oscillations have nearly flattened, indicating more significant energy dissipation for stronger nonlocality and more effective dissipation of vibrational energy in the system.
Figure 4: Effect of the horizon size $\Delta$ on the damping behavior of the channel height oscillations at an early time of $T=0.02$, showing the stronger wave damping as $\Delta$ increases. The simulation parameters are $St=10$, $\beta=10^{5}$, and $Re=10^{-5}$.